OpenStudy (anonymous):

a vehicle with a mass of 800kg experience a friction force of 180N and travels at a constant speed of 54km/h. Calculate the power required from the engine to travel at the constant speed.

OpenStudy (anonymous):

F1= 800kg times 9.8 m/s2 , 180 N Ff= 180N, opposing forces, calculate the difference F total; P= W/ t = Fxd / t= Ftotal X V ; change the unit of Km/ h to m/s standard units.

OpenStudy (anonymous):

for constant speed ,the vehicle should be in equilibrium...i.e we should apply 180N force for equilibrium . now, the force applied by engine=180N hence power applied by engine=applied force*speed of vehicle  calculate power..